OpenStudy (anonymous):

Solve x4 = 36x2. Identify any double roots A. {0 (double root), 6, –6} B. {–3, 3, 2 (double root)} C. {0, 3, 2} D. {6, –6}

OpenStudy (anonymous):

I put D as the answer because 6,-6 is the answer but I am wondering if A may be the answer.

OpenStudy (bahrom7893):

\[x^4-36x^2=0\]\[x^2(x^2-36)=0\]\[x*x(x-6)(x+6)=0\]

OpenStudy (anonymous):

This is what I got x^2(x+6)(x-6)

OpenStudy (anonymous):

since there is x^2 does that mean that 0 might be the double root

OpenStudy (bahrom7893):

Yes it does

OpenStudy (anonymous):

so the answer would be A right

OpenStudy (bahrom7893):

yup

OpenStudy (anonymous):

Thank you so much. I was really confused

OpenStudy (bahrom7893):

Np